Suppose that you bought a car in mid-2009 for $10,000. If its value (in current dollars) depreciates steadily at 12% per year, what will its value (in current dollars) be in mid 2012?
Answer by math_tutor2020(3817)


P = 10,000 = initial value
r = -12% = -0.12 = rate of growth (negative growth means depreciation)
t = 2012-2009 = 3 years

A = P*(1+r)^t
A = 10000*(1+(-0.12))^3
A = 10000*(1-0.12)^3
A = 10000*(0.88)^3
A = 6814.72

Answer: 6814.72 dollars